# RevealNet: Distributed Traffic Correlation for Attack Attribution on Programmable Networks

### v1

@@ 4.3 Flow Identification and Tracking
4.4 Compact Flow Features’ Representation and Correlation
@@
To minimize the amount of data that must be kept by a switch for each active flow, we explore existing methods of generating compact feature vectors. These approaches have been found to be applicable to traffic correlation workloads as well as other ML-based security tasks focused on flow analysis.
Traffic aggregation matrix. We first adopt a methodology (for generating feature vectors) which follows a traffic aggregation matrix (TAM) [39] approach. This matrix records metadata about packets transmitted per flow across multiple bins of seconds each, for a maximum of seconds, thereby storing a flow-level feature vector in a fixed-size data structure, consuming significantly less memory when compared to storing individual packet features. Similarly to Lopes et al. [25], we generate a single-row TAM per flow, where each TAM bin tracks the total number of packets transmitted by a flow within that bin’s time interval.
While storing complete TAM feature vectors in a switch’s data plane might be feasible, their memory footprint may compromise the concurrent storage of many flows simultaneously crossing the switch (see TAMs’ trade-offs in §5.3).
Flow sketching. To further compress TAM feature vectors while faithfully retaining flows’ characteristics, we use sketching techniques based on vector projection methods [12, 30]. Briefly, let the TAM feature vector for a flow be . Sketching algorithms transform into a lower-dimensional vector , where . The parameterization of such sketches enables us to trade-off the usage of switch memory (and thus, the total amount of concurrent flows that can be measured) with correlation accuracy.
We integrate these mathematical constructs into RevealNet’s P4 per-packet data plane processing logic, contrasting the use of two prominent sketching algorithms, proposed in the traffic correlation literature (§2.1), as the main driver of RevealNet’s attack attribution mechanisms. We describe them below, before detailing how flows’ sketches can be compared towards realizing flow correlation.
Coskun et al. [12]. This work proposes an online sketching method that first bins packets into discrete time slots (i.e., a packet count-based TAM vector) and then leverages linear transformations to generate a compact integer-array sketch representation of a flow. Sketches are computed on-the-fly without the need for temporarily storing the complete TAM feature vector. As the basis for these transformations, we use a random projection matrix whose entries are independently drawn from a Bernoulli distribution (i.e., each entry is either +1 or -1 with equal probability). This projection preserves the inherent structure of the packet-count vector and produces a sketch that contains only integer values, offering low per-packet overhead and robustness to network perturbations. The sketches can be binarized to save space and enable more efficient comparisons.
Nasr et al. [30]. This work proposes the aggregation of raw traffic features into a feature vector, which is then compressed using a sensing matrix into a lower-dimensional sketch. satisfies the restricted isometry property, allowing Euclidean distances between features to be preserved in the compressed domain.
In our implementation, we conducted two adaptations to Nasr et al.’s [30] original approach. First, since this scheme originally compresses the full feature vector, we implement a continuous update of a flow’s feature sketch every time a packet is processed, thus replicating the online sketching nature of Coskun et al. [12]. Second, since sensing matrices are instantiated as random Gaussian matrices with std. dev. , these contain floating-point entries which are not supported by P4 switches (§2.2). To address this issue, we scale the matrix by a constant factor—, in our implementation—to convert its entries to integer values without losing significant precision. Sketching is then performed on the P4 switch using this scaled matrix, enabling integer-only arithmetic.
Table 1 depicts the storage requirements for holding a TAM for a single flow, when contrasted to the storage required to hold the sketches we consider [12, 30].
| Method / Storage | Proj. Matrix | Flows | Total (bits) | Total (bits) w/ heur. (§4.5) |
| Nasr (integer sketch) | (integers) | (integers) | ||
| Coskun (integer sketch) | (integers) | (integers) | ||
| Coskun (binary sketch) | (integers) | (bits) | ||
@@
Correlation. The final step in RevealNet’ attack attribution pipeline involves correlating feature vectors to identify whether two flows collected at different vantage points originate from the same source. In RevealNet, correlation is based on computing a statistical distance or similarity between the sketches. Different sketching methods use disparate metrics for enacting said comparisons. Coskun et al. [12] use the Hamming distance to compare sketches while Nasr et al. [30] employ cosine similarity. Our implementation relies on the same metrics.
4.5 Heuristic Optimizations for Attack Attribution
@@
Packet count heuristic. Flows with a total packet count similar to that of the malicious flow are more likely to be true matches. By bounding the acceptable packet count range using thresholds and , derived from the target malicious flow, we restrict correlation to flows with comparable traffic volumes.
We apply the heuristics one after the other. This two-step strategy reduces correlation complexity from the baseline —where is the total number of outgoing flows observed in (see §3) and the cost of a single comparison between two flows’ feature vectors—to . Here, and represent the reduced flow sets after the cumulative timestamp and packet count filtering, respectively. Since flows are pre-sorted by timestamp, identifying requires only via binary search. Filtering by packet count is linear in , yielding in . We then perform flows’ feature vector comparisons only on this set, which incurs a cost of .
To implement the heuristics, the data plane of each switch maintains two separate tables with a number of rows equal to the number of flows. Each entry stores auxiliary metadata: 48 bits for a flow’s creation timestamp and 32 bits for that flow’s total packet count. This results in a storage overhead of , where denotes the number of flows observed by a switch (see Table 1).
4.6 Implementation
We built a prototype of RevealNet using bmv2 [35], the reference P4 software switch. The data plane logic, including flow identification and sketching operations (for either sketch), was implemented in 500 lines of P4_16 code. In turn, RevealNet’s control plane logic was written in 300 lines of Python code. This includes the installation of tables and rules supporting flow identification and sketching operations in the data plane, as well as fetching flows’ feature vectors via reads to data plane registers for enabling the correlation backbone.
5 Evaluation
5.1 Evaluation Goals and Metrics
We wish to evaluate the practicality of RevealNet along three key dimensions:
@@ 5.1 Evaluation Goals and Metrics
Efficiency. We quantify the computational cost associated with flow correlation via the number of pairwise comparisons between detected malicious flows and the outgoing flows observed by cooperating networks. The computational effort across all switches can be expressed as the sum of the pairwise comparisons for each switch: , where represents the number of malicious flows detected and represents the number of outgoing flows observed by switch .
@@
5.2 Evaluation Methodology
We now describe the datasets used in our evaluation, followed by details on the parameterization of RevealNet’s data structures and heuristics.
| Dataset | Category | Description | Flows | Duration (s) | |
@@ 5.2 Evaluation Methodology
| Adload | Adware | Static resources for PPI adware. | 125 417 | 602 | 54.80 |
@@
Datasets used for attack attribution. We use six labelled network traffic datasets, released by Fu et al. [15], as a target of RevealNet’ attack attribution capabilities. These datasets were compiled from a combination of Fu et al.’s own experimental data and traffic traces from the WIDE MAWI project in Tokyo, Japan. For exercising RevealNet’s generalizability, we selected each dataset from six different categories of attacks collected by Fu et al. (see Table 2 for an overview of each dataset). Each of the datasets contains a different type of network attack along with background benign traffic. Fu et al.’s data records detail per-packet five-tuples in .csv files, along with packets’ timestamps and labels (benign/malicious), allowing us to carve out individual flows identified by these 5-tuples. Each dataset accounts for more than 100k flows, and the ratio of benign to malicious traffic is at least 39:1 (Dridex) and at most 1312:1 (Oracle).
For simplicity, we assume that the network IDS deployed within each RevealNet-enabled network acts as an oracle that can perfectly distinguish between benign and malicious flows. While this is within the realm of practicality for the datasets we considered [15], we recall that our goal is not to perform accurate malware classification, but rather to act on the alerts produced by IDSes (§3).
Emulating vantage points and network conditions. Since the above datasets were collected at a single network vantage point and do not include raw packet traces that can be transparently replayed across some network topology (real or emulated) by special-purpose software such as tcpreplay, they cannot be directly used for correlation experiments across different networks, as required by RevealNet. To tackle this issue, and similarly to [12], we simulate the acquisition of two separate observations for each flow at different vantage points within RevealNet-enabled networks: a) at the border router of a cooperating network where hosts originate benign/malicious traffic, and; b) at the border router of a network which is targeted by some attack. We also assume that all flows in each of Fu et al. [15]’s datasets originate from a cooperating network and traverse (or target devices within) the attacked network. To facilitate this setup, we implemented a simulator that models WAN traffic being relayed via an intermediate proxy node. We produce a second set of datasets that represent traffic observations across the WAN, where packets experience an average latency increase of 200ms between any two vantage points. The simulator also supports the injection of perturbations, e.g., packet loss, allowing us to assess the robustness of flow correlation under degraded network conditions.
Parameterization of RevealNet’s sketches. Each dataset from Fu et al. [15] spans 45–65s of traffic. To explore the impact of temporal granularity in flow feature collection, we generated TAM time bins () of 0.1s, 0.5s, and 1s. We performed preliminary experiments using different sketch lengths (), keeping as a baseline. A sketch length of 5 slightly improved TPR by up to +1.52% but at the cost of a substantial increase in FPR, reaching +114.84%. Conversely, using provided no consistent TPR gains and resulted in mixed FPR outcomes (ranging from –20.4% to +92.65%), along with added storage overhead. Overall, for all datasets we considered, strikes a favourable balance between accuracy and efficiency (see §5.3). Still, as previously studied [12], this parameter may need to be tuned for flows with different traits.
We follow the original methodology of each sketch to compute correlation scores. For Coskun et al.’s sketch, we use Hamming distance and consider a match to be a true positive only when the distance between sketches is 0. For Nasr et al.’s sketch, we use cosine similarity, requiring a score of 1 for an exact match. We evaluate correlation using TAMs with both Hamming distance and cosine similarity, applying the same thresholds to define true positives. We adopt these thresholds to reflect high-confidence matches in attack attribution, where false associations can be especially harmful to benign users. Indeed, we experimented with relaxed correlation thresholds across various time bins , but found that these looser criteria offered only marginal improvements in true positive rates while significantly increasing false positives. For instance, for Nasr et al.’s sketches’ setups, a cosine similarity threshold of 0.9 yielded modest increases in TPR (up to 2.54%) but introduced significantly higher FPR (up to 373.04%).
@@
Configuration of RevealNet’s heuristics. Network topology and flows’ characteristics can significantly influence the choice of RevealNet’s heuristics’ configurations [6]. However, since the datasets we use to evaluate RevealNet (see Table 2) share similar traits on flow durations, we configure our heuristics to be consistent across all datasets. For the timing-based and packet count heuristics, we empirically found that a window of seconds and a threshold of traffic volume filters out irrelevant candidate flows while retaining most valid matches. As described in §5.3, applying these heuristics to the bitcoinminer dataset helps reduce the number of comparisons per attacking flow by 3 orders of magnitude, without substantially sacrificing correlation accuracy.
5.3 Evaluation Results
@@
We now describe the main results obtained during our evaluation of RevealNet. Throughout this section, we centre on the bitcoinminer dataset since it: a) has a representative benign-to-malicious flow ratio of 625:1, close to the median of all datasets; b) exhibits the second-largest temporal span, and; c) has a relatively smaller number of flows, facilitating faster experimentation. Results for other datasets, which exhibit similar trends, are deferred to Appendix 0.A.1.
| Dataset | None | Creation Time | Packet Count | Both |
| Dridex | 128 626 | 19 409 | 2 403 | 824 |
| Adload | 126 019 | 17 386 | 1 168 | 258 |
| Snojan | 208 330 | 20 975 | 2 735 | 693 |
| Oracle | 294 334 | 41 580 | 95 | 84 |
| Bitcoinminer | 125 620 | 17 288 | 791 | 207 |
@@
| Method/ Storage | Per flow (in Bytes) | Stored Flows (in 256 MB) |
| Nasr et al. (integer) | 40 | |
| Coskun et al. (integer) | 40 | |
| Coskun et al. (binary) | 1.25 | |
| TAM (0.1s bins) | 2864 | |
| TAM (0.5s bins) | 576 | |
@@
Compact sketches obtain correlation accuracy equivalent to TAMs. Figure 5 shows the TPR and FPR achieved by different parameterizations of the TAM and sketching approaches considered in our work, when correlating flows pertaining to the bitcoinminer dataset. The figure suggests that sketches attain a comparable correlation performance vs. TAMs, making them highly attractive due to their smaller memory overhead. Focusing on the results obtained without the use of heuristics (solid bars), the integer sketching method from Coskun et al. [12] achieves equivalent accuracy (TPR: , FPR: ) to TAM (in 0.1 seconds time bin and Hamming distance) while offering significant space savings—indeed, Figure 6 illustrates that TAM’s memory footprint can be up to larger than that of sketches under finer-grained binnings (e.g., ).
To assess whether the benefits of sketches are still applicable under perturbed network conditions, we performed additional experiments where we simulated the introduction of random packet drops at rates of –. We now present our findings assuming that heuristics were deployed. As shown in Figure 5, both TPR and FPR declined under network perturbations (bars with markers), consistent with prior observations [12, 30]. For instance, the integer sketch (0.1s) maintained a TPR of and FPR of , equivalent to that of TAM under the packet drop of 5%. Interestingly, the binary sketching variant yielded a significantly higher TPR () compared to both integer sketches and TAM in the same conditions, albeit with a modest increase in FPR.
Overall, the above results suggest that, even in noisy environments, sketches preserve the detection characteristics of more resource-intensive TAM configurations and can be relied upon for realizing RevealNet’s correlation backbone. While not the main focus of our paper, we reckon that advanced traffic shaping techniques—beyond usual network perturbations—are challenging for accurate flow correlation, making this process significantly harder, if not infeasible [32, 25].
Heuristics reduce complexity and foster improved correlation. Table 4 illustrates the impact of heuristics on reducing the flow comparisons performed by RevealNet. Without heuristics, comparisons range from 125k (bitcoinminer) to nearly 295k (penetho and oracle). With heuristics applied, this drops to 207, 454, and 84, respectively, thus reducing the computational effort involved in the correlation workload by at least three orders of magnitude.
@@
Sketches allow for storing more flows concurrently. Table 4 presents the approximated number of flows that can be stored in a Tofino v1 P4 switch equipped with 256 MB of SRAM [54], for the various feature extraction methods under analysis. All sketches are configured with a length of . For Coskun et al. and Nasr et al., the sketching process requires storing a projection matrix of size (§4.4), which introduces a storage overhead of 24 400B, 4 920B, and 2 480B for TAMs based on 0.1s, 0.5s, and 1s bins (), respectively.
Although sketches require this fixed overhead, they dramatically improve storage capacity. For example, Coskun et al.’s binary sketch supports storing up to flows, compared to just with TAM at s granularity—the most memory-intensive setting in our study. Other sketching methods show similar scalability, reinforcing that sketch-based correlation is well-suited for memory-constrained P4 switches that must handle large flow volumes.
| Method | Centralized | Distributed (RevealNet) | OH Red. (%) | ||||||
| Swc CS | Swa CS | Total | Sa CM | CM Swc | Swc CM | Total | |||
| Coskun et al. (binary) | 1 193 390 | 2 020 | 1 195 410 | 2 020 | 38 380 | 736 896 | 777 296 | 35.0% | |
| Coskun et al. (integer) | 38 188 480 | 64 640 | 38 253 120 | 64 640 | 1 227 680 | 736 896 | 2 029 216 | 94.7% | |
| Nasr et al. (integer) | 38 188 480 | 64 640 | 38 253 120 | 64 640 | 1 227 680 | 736 896 | 2 029 216 | 94.7% | |
Distributed correlation saves bandwidth. So far, our evaluation considers a single-switch setup. In practice, however, attack attribution spans multiple P4 switches located across different cooperating networks. In this distributed setting, correlation scales naturally in an “embarrassingly parallel” fashion: each switch handles its local traffic and performs correlation independently. We now gauge the communication overheads imposed by RevealNet, comparing them to traditional centralized attack attribution deployments.
@@
To evaluate the communication overhead of centralized vs. distributed correlation, we simulate a topology with 20 RevealNet switches: 19 monitoring outgoing flows at cooperating networks (Swc), and one observing incoming flows at an attacked network (Swa). Assuming an even distribution of flows sourced from bitcoinminer (where we assume all flows to originate in cooperating networks and traverse the attacked network), each Swc sees 6 281 outgoing flows, while Swa sees a total of 119 339 incoming flows, out of which 202 are malicious. In a centralized setup, each Sw sends all observed flows’ feature vectors to a CS, while the Swa sends its 202 feature vectors. In RevealNet, Swa sends the 202 feature vectors to the CM, which relays them to all Swc. Each Swc performs correlation locally and returns 202 matched flow tuples (192 bits each) to CM.
Table 5 shows a breakdown of the total communication involved in both scenarios. In our example setup with 20 switches/networks, RevealNet’s distributed design reduces bandwidth usage by 35%–94.7%, depending on the used sketch.
6 Conclusions
In this work, we introduced RevealNet, a practical framework for distributed attack attribution across multiple cooperating networks. By leveraging compact sketch-based data structures and the orchestration of programmable network elements, RevealNet is able to accurately correlate malicious flows while maintaining low computational and communication overheads. Our evaluation across multiple datasets suggests that flow correlation can be effectively pushed into the network fabric itself, paving the way for scalable attack attribution systems.
Appendix 0.A Appendix
0.A.1 Correlation Effectiveness across Multiple Datasets
Figures 11–11 show RevealNet’s correlation effectiveness on additional datasets from Fu et al. [15], exhibiting similar trends to those observed in §5.

### v2

@@ IV-C Flow Identification and Tracking
IV-D Compact Flow Features’ Representation and Correlation
@@
To minimize the amount of data that must be kept by a switch for each active flow, we explore existing methods of generating compact feature vectors. These have been found to be applicable to traffic correlation workloads as well as other ML-based security tasks focused on flow analysis.
Traffic aggregation matrix. We first adopt a methodology (for generating feature vectors) which builds a traffic aggregation matrix (TAM) [18]. This matrix records metadata about packets transmitted per flow across multiple bins of seconds each, for a maximum of seconds, thereby storing a flow-level feature vector in a fixed-size data structure, consuming significantly less memory when compared to storing individual packet features. Like Lopes et al. [5], we generate a single-row TAM per flow, where each TAM bin tracks the number of packets transmitted by a flow within that bin’s time interval.
While storing complete TAM feature vectors in a switch’s data plane might be feasible, their memory footprint may compromise the concurrent storage of many flows simultaneously crossing the switch (see TAMs’ trade-offs in §V-C).
| Method / Storage | Proj. Matrix | Flows | Total (bits) | Total (bits) w/ heur. (§IV-E) |
| Nasr et al. (integer sketch) | (integers) | (integers) | ||
| Coskun et al. (integer sketch) | (integers) | (integers) | ||
| Coskun et al. (binary sketch) | (integers) | (bits) | ||
@@
Flow sketching. To further compress TAM feature vectors while retaining flows’ characteristics, we use sketching techniques based on vector projection methods [11, 9]. Briefly, let the TAM feature vector for a flow be . Sketching algorithms transform into a lower-dimensional vector , where . The parameterization of such sketches enables us to trade-off the usage of switch memory (and thus, the number of concurrent flows that can be measured) with correlation accuracy.
We integrate these constructs into RevealNet’s data plane processing logic, contrasting the use of two prominent sketching algorithms, proposed in the traffic correlation literature (§II-A), as the main driver of RevealNet’s attack attribution mechanisms. We describe them below, and detail how flows’ sketches can be compared towards realizing flow correlation.
Coskun et al. [11] propose an online sketching method that first bins packets into discrete time slots (i.e., a packet count-based TAM vector) and then leverages linear transformations to generate a compact integer-array sketch representation of a flow. Sketches are computed on-the-fly without the need for temporarily storing the complete TAM feature vector. As the basis for these transformations, we use a random projection matrix whose entries are independently drawn from a Bernoulli distribution (i.e., each entry is either +1 or -1 with equal probability). This projection preserves the structure of the packet-count vector and produces a sketch that contains only integer values, offering low per-packet overhead and robustness to network perturbations. The sketches can be binarized to save space and enable more efficient comparisons.
Nasr et al. [9] propose the aggregation of raw traffic features into a feature vector, which is then compressed using a sensing matrix into a lower-dimensional sketch. satisfies the restricted isometry property, allowing Euclidean distances between features to be preserved in the compressed domain.
In our implementation, we conducted two adaptations to Nasr et al.’s [9] original approach. First, since this scheme originally compresses the full feature vector, we implement a continuous update of a flow’s feature sketch every time a packet is processed, thus replicating the online sketching nature of Coskun et al. [11]. Second, since sensing matrices are instantiated as random Gaussian matrices with std. dev. , these contain floating-point entries which are not supported by P4 switches (§II-B). To address this issue, we scale the matrix by a constant factor—, in our implementation—to convert its entries to integer values without losing significant precision. Sketching is then performed on the P4 switch using this scaled matrix, enabling integer-only arithmetic.
Table I depicts the storage requirements for holding a TAM for a single flow, when contrasted to the storage required to hold the sketches we consider [11, 9].
Correlation. The final step in RevealNet’ pipeline involves correlating feature vectors to identify whether two flows collected at different vantage points originate from the same source. In RevealNet, correlation is based on computing a statistical distance or similarity between the sketches. Different sketching methods use disparate metrics for enacting said comparisons. Coskun et al. [11] use the Hamming distance to compare sketches while Nasr et al. [9] employ cosine similarity. Our implementation relies on the same metrics.
IV-E Heuristic Optimizations for Attack Attribution
@@
Packet count heuristic. Flows with packet counts akin to that of the malicious flow are more likely to be true matches. By bounding the acceptable packet count range using thresholds and , derived from the target malicious flow, we restrict correlation to flows with comparable traffic volumes.
We apply the heuristics one after the other. This two-step strategy reduces correlation complexity from the baseline —where is the total number of outgoing flows observed in (see §III) and the cost of a single comparison between two flows’ feature vectors—to . Here, and represent the reduced flow sets after the cumulative timestamp and packet count filtering, respectively. Since flows are pre-sorted by timestamp, identifying requires only via binary search. Filtering by packet count is linear in , yielding in . We then perform flows’ feature vector comparisons only on this set, which incurs a cost of .
To implement the heuristics, the data plane of each switch maintains two separate tables with a number of rows equal to the number of flows. Each entry stores auxiliary metadata: 48 bits for a flow’s creation timestamp and 32 bits for that flow’s total packet count. This results in a storage overhead of , where denotes the number of flows observed by a switch (see Table I).
IV-F Implementation
We built a prototype of RevealNet using bmv2, the reference P4 software switch. The data plane logic, including flow identification and sketching operations (for either sketch), was implemented in 500 lines of P416 code. In turn, RevealNet’s control plane logic was written in 300 lines of Python code. This includes the installation of tables and rules supporting flow identification and sketching operations in the data plane, as well as fetching flows’ feature vectors via reads to data plane registers for enabling the correlation backbone.
V Evaluation
V-A Evaluation Goals and Metrics
We assess RevealNet’s practicality along three facets:
@@ V-A Evaluation Goals and Metrics
Efficiency. We quantify the computational cost linked to flow correlation via the number of pairwise comparisons between detected malicious flows and the outgoing flows observed by cooperating networks. The computational effort across all switches is expressed as the sum of the pairwise comparisons for each switch: , where represents the number of malicious flows detected and represents the number of outgoing flows observed by switch .
@@
V-B Evaluation Methodology
We now describe the datasets used in our evaluation, and the configuration of RevealNet’s data structures and heuristics.
Datasets used for attack attribution. We use six labelled network traffic datasets, released by Fu et al. [30], as a target of RevealNet’ attack attribution capabilities. These datasets were compiled from a combination of Fu et al.’s own experimental data and traffic traces from the WIDE MAWI project in Tokyo, Japan. For exercising RevealNet’s generalizability, we selected each dataset from six different categories of attacks collected by Fu et al. (see Table II for a description). Each dataset contains a different type of network attack along with background benign traffic. Fu et al.’s data details per-packet five-tuples in .csv files, along with packets’ timestamps and labels (benign/malicious), allowing us to carve out individual flows identified by these 5-tuples. Each dataset accounts for more than 100k flows, and the ratio of benign to malicious traffic is at least 39:1 (Dridex) and at most 1312:1 (Oracle).
For simplicity, we assume that the network IDS deployed within each RevealNet-enabled network acts as an oracle that can perfectly distinguish between benign and malicious flows. While this assumption is already aligned with the capabilities of state-of-the-art IDSes for the datasets we considered [30], we recall that our goal is not to perform accurate malware classification, but rather to act on IDSes’ alerts (§III). In practice, false positives would increase the number of correlation tasks to be performed, while false negatives would prevent the attribution of some attacks, since those malicious flows would never be sent to the correlation manager.
Simulating vantage points and network conditions. Since the above datasets were collected at a single network vantage point and do not include raw packet traces that can be transparently replayed across some network topology (real or emulated) by special-purpose software such as tcpreplay, they cannot be directly used for correlation experiments across different networks, as required by RevealNet.
To tackle this issue, akin to [11], we simulate the acquisition of two separate observations for each flow at different vantage points within RevealNet-enabled networks: a) at the border router of a cooperating network where hosts originate benign/malicious traffic, and; b) at the border router of a network which is targeted by some attack. We also assume that all flows in each dataset originate from a cooperating network and traverse (or target devices within) the attacked network.
To facilitate this setup, we implemented a simulator that models WAN traffic relayed via a proxy node. We used the simulator to augment the traffic traces of Fu et al. [30], reproducing the observation of flows across two vantage points to mimic distributed monitoring. The resulting traces, which we use throughout our evaluation, capture traffic across the WAN where packets incur an average latency increase of 200ms between any two vantage points; this falls within the range of end-to-end latencies between geo-distributed client-proxy-server nodes in stepping-stone scenarios that consider wide-area paths [31]. The simulator also supports the injection of packet losses for us to assess the robustness of flow correlation under network perturbations.
Parameterization of RevealNet’s sketches. Each dataset from Fu et al. [30] spans 45–65s of traffic. To explore the impact of temporal granularity in flow feature collection, we generated TAM time bins () of 0.1s, 0.5s, and 1s. We performed preliminary experiments using different sketch lengths (), keeping as a baseline. A sketch length of 5 slightly improved TPR by up to +1.52% but at the cost of a substantial increase in FPR, reaching +114.84%. Conversely, using provided no consistent TPR gains and resulted in mixed FPR outcomes (ranging from –20.4% to +92.65%), along with added storage overhead. Overall, for all datasets we considered, strikes a favourable balance between accuracy and efficiency (§V-C). Still, this parameter may need to be tuned for flows with different traits [11].
We follow the original methodology of each sketch to compute correlation scores. For Coskun et al.’s sketch, we use Hamming distance and consider a match to be a true positive only when the distance between sketches is 0. For Nasr et al.’s sketch, we use cosine similarity, requiring a score of 1 for an exact match. We evaluate TAMs with both Hamming distance and cosine similarity, applying the same thresholds to define true positives. We adopt these thresholds to reflect high-confidence matches in attack attribution, where false associations can be harmful to benign users. Indeed, we experimented with relaxed correlation thresholds across various time bins , but found that these looser criteria offered only marginal improvements in TPR while significantly increasing FPR. For instance, for Nasr et al.’s sketches, a cosine similarity threshold of 0.9 yielded modest increases in TPR (up to 2.54%) but significantly higher FPR (up to 373.04%).
| Dataset | Category | Description | Flows | Span (s) | |
@@ V-B Evaluation Methodology
| Adload | Adware | Resources for PPI adware. | 125 417 | 602 | 54.80 |
@@
Configuration of RevealNet’s heuristics. Network topology and flows’ characteristics can influence RevealNet’s heuristics’ configurations [29]. However, since the datasets we consider (see Table II) share similar traits on flow durations, we configure our heuristics to be consistent across all datasets. For the timing-based and packet count heuristics, we empirically found that a window of seconds and a threshold of traffic volume filters out irrelevant candidate flows while retaining most valid matches. As described in §V-C, applying these heuristics to bitcoinminer reduces the number of comparisons per attacking flow by 3 orders of magnitude, without substantially sacrificing correlation accuracy.
V-C Evaluation Results
@@
We now describe the main results obtained during our evaluation of RevealNet. We centre our exposition on the bitcoinminer dataset since it: a) has a representative benign-to-malicious flow ratio of 625:1, close to the median of all datasets; b) exhibits the second-largest temporal span, and; c) has a relatively smaller number of flows, facilitating faster experimentation. Results for other datasets exhibit similar trends, and we defer their discussion to Appendix A.
Compact sketches obtain correlation accuracy equivalent to TAMs. Figure 4 shows the TPR and FPR achieved by different parameterizations of the TAM and sketching approaches considered in our work, when correlating flows pertaining to the bitcoinminer dataset. The figure suggests that sketches attain a comparable correlation performance vs. TAMs, making them highly attractive due to their smaller memory overhead. Focusing on the results obtained without the use of heuristics (solid bars), the integer sketching method from Coskun et al. [11] achieves equivalent accuracy (TPR: , FPR: ) to TAM (in 0.1 seconds time bin and Hamming distance) while offering significant space savings—indeed, Figure 5 illustrates that TAM’s memory footprint can be up to larger than that of sketches under finer-grained binnings (e.g., ) across all datasets.
| Dataset | None |
Both | ||||
| Dridex | 128 626 | 19 409 | 2 403 | 824 | ||||
| Adload | 126 019 | 17 386 | 1 168 | 258 | ||||
| Snojan | 208 330 | 20 975 | 2 735 | 693 | ||||
| Oracle | 294 334 | 41 580 | 95 | 84 | ||||
| Bitcoinminer | 125 620 | 17 288 | 791 | 207 | ||||
@@
| Method/ Storage |
| Coskun et al. (bin.) | 1.25 | |||||
| Coskun et al. (int.) | 40 | |||||
| Nasr et al. (int.) | 40 | |||||
| TAM (0.1s bins) | 2864 | |||||
| TAM (0.5s bins) | 576 | |||||
@@
| Method | Centralized | Distributed (RevealNet) | OH Red. (%) | ||||||
| Swc CS | Swa CS | Total | Swa CM | CM Swc | Swc CM | Total | |||
| Coskun et al. (bin.) | 1 193 390 | 2 020 | 1 195 410 | 2 020 | 38 380 | 736 896 | 777 296 | 35.0% | |
| Coskun et al. (int.) | 38 188 480 | 64 640 | 38 253 120 | 64 640 | 1 227 680 | 736 896 | 2 029 216 | 94.7% | |
| Nasr et al. (int.) | 38 188 480 | 64 640 | 38 253 120 | 64 640 | 1 227 680 | 736 896 | 2 029 216 | 94.7% | |
To assess if sketches can succeed under perturbed network conditions, we simulated random packet drops at – rates. We now present our findings assuming that heuristics were deployed. As shown in Figure 6, both TPR and FPR declined under packet loss (bars with markers), consistent with prior observations [11, 9]. For instance, the integer sketch (0.1s) maintained a TPR of and FPR of , equivalent to that of TAM under the packet drop of 5%. The binary sketching variant yielded a significantly higher TPR () compared to both integer sketches and TAM in the same conditions, albeit with a modest increase in FPR.
Overall, the above results suggest that, even in noisy environments, sketches preserve the detection characteristics of more resource-intensive TAM configurations and can be relied upon for realizing RevealNet’s correlation backbone. Yet, advanced traffic shaping techniques beyond network perturbations, such as traffic shaping proxies [32], present an inherent limitation to traffic correlation schemes, making this process significantly harder, if not infeasible [8, 5].
Heuristics reduce complexity and foster improved correlation. Table IV illustrates the impact of heuristics on reducing the flow comparisons performed by RevealNet. Without heuristics, comparisons range from 125k (bitcoinminer) to nearly 295k (penetho and oracle). With heuristics applied, this drops to 207, 454, and 84, respectively, thus reducing the computational effort involved in the correlation workload by at least three orders of magnitude.
@@
Sketches allow for storing more flows concurrently. Table IV presents the approximated number of flows that can be stored in a Tofino v1 P4 switch equipped with 256 MB of SRAM, for the various feature extraction methods under analysis. All sketches are configured with a length of . For Coskun et al. and Nasr et al., the sketching process requires storing a projection matrix of size (§IV-D), which introduces a storage overhead of 24 400B, 4 920B, and 2 480B for TAMs based on 0.1s, 0.5s, and 1s bins (), respectively.
Though sketches require this fixed overhead, they dramatically improve storage capacity. Coskun et al.’s binary sketch stores up to flows, compared to just with TAM at s granularity—our most memory-intensive setting. Other sketches show similar scalability, reinforcing that sketch-based correlation is well-suited for memory-constrained P4 switches that must handle large flow volumes.
Distributed correlation saves bandwidth. Attack attribution spans multiple P4 switches distributed across different cooperating networks. Thus, correlation scales in an “embarrassingly parallel” fashion: each switch handles its local traffic and performs correlation independently. We now gauge the communication overheads imposed by RevealNet, comparing them to those of centralized attack attribution deployments.
@@
To gauge the communication overhead of centralized vs. distributed correlation, we simulate a topology with 20 RevealNet switches: 19 monitoring outgoing flows at cooperating networks (Swc), and one observing incoming flows at an attacked network (Swa). Assuming an even distribution of flows sourced from bitcoinminer (where all flows originate in cooperating networks and traverse the attacked network), each Swc sees 6 281 outgoing flows, while Swa sees a total of 119 339 incoming flows, out of which 202 are malicious. In a centralized setup, each Sw sends all observed flows’ feature vectors to a CS, while the Swa sends its 202 feature vectors. In RevealNet, Swa sends the 202 feature vectors to the CM, which relays them to all Swc. Each Swc performs correlation locally and returns 202 matched flow tuples (192 bits each) to the CM. Table V shows the communication involved in both scenarios. RevealNet’s distributed design reduces bandwidth usage by 35%–94.7%, depending on the sketch.
VI Conclusion
We introduced RevealNet, a practical framework for distributed attack attribution across cooperating networks. By using compact sketch-based data structures and the orchestration of programmable network elements, RevealNet is able to accurately correlate malicious flows while maintaining low computational and communication overheads. Our evaluation suggests that flow correlation can be effectively pushed into the network fabric, paving the way for scalable attack attribution.
Appendix A Appendix
A-A Correlation Effectiveness across Multiple Datasets
Figures 7(a)–7(e) show RevealNet’s correlation effectiveness on additional datasets from Fu et al.[30], showing similar trends to those discussed in Section V on bitcoinminer.
The TPR and FPR obtained across all datasets indicate that using the integer sketch by Coskun [11] achieves performance similar to TAM, while requiring 10% to 60% less memory (see Figure 5). This suggests that sketching can be effective for reducing memory usage. The results also show that the binary sketch by Coskun et al. [11] achieves higher TPR than TAM in some cases, but with a higher FPR. Both sketches by Nasr et al. [9] perform either similarly to or worse than Coskun’s sketches. These findings further support the claim made in Section V that sketches can be used to reduce memory and computation costs while maintaining comparable performance.